# Bandit and Delayed Feedback in Online Structured Prediction

### v1

@@ Abstract
Online structured prediction is a task of sequentially predicting outputs with complex structures based on inputs and past observations, encompassing online classification. Recent studies showed that in the full information setup, we can achieve finite bounds on the surrogate regret, i.e., the extra target loss relative to the best possible surrogate loss. In practice, however, full information feedback is often unrealistic as it requires immediate access to the whole structure of complex outputs. Motivated by this, we propose algorithms that work with less demanding feedback, bandit and delayed feedback. For the bandit setting, using a standard inverse-weighted gradient estimator, we achieve a surrogate regret bound of $O(\sqrt{KT})$ for the time horizon $T$ and the size of the output set $K$. However, $K$ can be extremely large when outputs are highly complex, making this result less desirable. To address this, we propose an algorithm that achieves a surrogate regret bound of $O(T^{2/3})$, which is independent of $K$. This is enabled with a carefully designed pseudo-inverse matrix estimator. Furthermore, for the delayed full information feedback setup, we obtain a surrogate regret bound of $O(D^{2/3}T^{1/3})$ for the delay time $D$. We also provide algorithms for the delayed bandit feedback setup. Finally, we numerically evaluate the performance of the proposed algorithms in online classification with bandit feedback.
@@ 1 Introduction
In many machine learning problems, given an input vector from a vector space $\mathcal{X}$, we aim to predict a vector in a finite output space $\mathcal{Y}$. Multiclass classification is one of the simplest examples, while in other cases output spaces may have more complex structures. Structured prediction refers to such a class of problems with structured output spaces, including multiclass classification, multilabel classification, ranking, and ordinal regression, and it has applications in various fields, ranging from natural language processing to bioinformatics (Tsochantaridis et al., 2005; BakIr et al., 2007). In structured prediction, training models that directly predict outputs in complex discrete output spaces is typically challenging. Therefore, we often adopt the surrogate loss framework (Bartlett et al., 2006)—define an intermediate space of score vectors and train models that estimate score vectors from inputs based on surrogate loss functions. Examples of surrogate losses include squared, logistic, and hinge losses, and a general framework encompassing them is the Fenchel–Young loss (Blondel et al., 2020), which we rely on in this study.
Structured prediction can be naturally extended to the online setting (Sakaue et al., 2024). In this online setting, at each round $t=1,\dots,T$, the environment selects an input-output pair $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{X}\times\mathcal{Y}$. A learner then predicts $\widehat{\bm{y}}_{t}\in\mathcal{Y}$ based on the input $\bm{x}_{t}$ and incurs a loss $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$, where $L:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is the target loss function. As with Sakaue et al. (2024), we focus on the simple yet fundamental case where the learner’s model for estimating score vectors is linear.
Table 1: Surrogate regret upper and lower bounds in online multiclass classification and online structured prediction. Here, $T$ is the time horizon, $K$ is the size of the set, $\mathcal{Y}$, of output vectors, and $D$ is the fixed delayed time. “FI” is the abbreviation of full information. Delayed feedback is considered only when “Delayed” appears in the feedback column. In the target loss column, “SELF*” means SELF that satisfies Assumption 3.7. Note that the $O(T^{2/3})$ bounds for SELF* in lines 6 and 9 do not explicitly depend on $K$ but on $d$; in the case of multiclass classification with the 0-1 loss, the dependence on $K$ appears as $d=K$.
Problem setup Feedback Target loss Surrogate regret bound
Van der Hoeven et al. (2021, Cor. 1) Binary classification Bandit 0-1 loss $\Omega(\sqrt{T})$ ($d=2$)
Van der Hoeven (2020, Thm. 4) Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
Van der Hoeven et al. (2021, Thm. 1) Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
Sakaue et al. (2024, Thms. 7 and 8) Structured prediction FI SELF $O(1)$
This work (Theorems 3.5 and 3.6) Structured prediction Bandit SELF $O(\sqrt{KT})$
@@
This work (Theorems 4.3 and 4.4) Structured prediction FI & Delayed SELF $O(D^{2/3}T^{1/3})$
This work (Theorem 5.1) Structured prediction Bandit & Delayed SELF $O(\sqrt{DKT})$
@@
The goal of the learner is to minimize the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$. On the other hand, the best the learner can do in the surrogate loss framework is to minimize the cumulative surrogate loss, namely $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$, where $\bm{U}:\mathcal{X}\to\mathbb{R}^{d}$ is the best offline linear estimator and $S:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a surrogate loss, which measures the discrepancy between the score vector $\bm{U}\bm{x}_{t}\in\mathbb{R}^{d}$ and $\bm{y}_{t}\in\mathcal{Y}$. Given this, a natural performance measure of the learner’s predictions is the following surrogate regret $\mathcal{R}_{T}$:
$\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}.$ (1)
The surrogate regret was introduced in the seminal paper by Van der Hoeven (2020) in the context of online multiclass classification. Recently, Sakaue et al. (2024) showed that a finite surrogate regret bound can be achieved for online structured prediction under full information feedback, i.e., the learner can observe $\bm{y}_{t}$ at the end of each round $t$.
However, the assumption that full information feedback is available is often demanding, especially when outputs have complex structures. For example, in sequential ad assortment on an advertising platform, we may be able to observe only the click-through rate but not which ads were clicked, which boils down to the bandit feedback setting (Kakade et al., 2008; Gentile & Orabona, 2014a). Also, we may only have access to feedback from a while ago when designing an ad assortment for a new user—namely, delayed feedback (Weinberger & Ordentlich, 2002; Manwani & Agarwal, 2023). Similar situations have led to a plethora of studies in various online learning settings. In combinatorial bandits, algorithms under bandit feedback (referred to as full-bandit feedback in their context), instead of full information feedback, have been widely studied (Cesa-Bianchi & Lugosi, 2012; Combes et al., 2015; Rejwan & Mansour, 2020; Du et al., 2021). Delayed feedback is also explored in various settings, including full information and bandit feedback (Joulani et al., 2013; Cesa-Bianchi et al., 2016). Due to space limitations, we defer a further discussion of the background to Appendix B.
@@
To extend the applicability of online structured prediction, this study develops online structured prediction algorithms that can handle weaker feedback—bandit feedback and delayed feedback—instead of full information feedback. As with Sakaue et al. (2024), we consider the case where target loss functions belong to a class called the Structured Encoding Loss Function (SELF) (Ciliberto et al., 2016; Blondel, 2019), a general class including the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking (see Section 2.4 for the definition). Table 1 summarizes the surrogate regret bounds provided in this study and comparisons with existing results.
@@
One of the challenges of bandit feedback is that the true output $\bm{y}_{t}$ is not observed, making it impossible to compute the true gradient of the surrogate loss. To deal with this, we use an inverse-weighted gradient estimator, a typical approach that assigns weights to gradients by the inverse of choosing each output, establishing an $O(\sqrt{KT})$ surrogate regret upper bounds (Theorems 3.5 and 3.6), where $K=\lvert\mathcal{Y}\rvert$ is the cardinality of $\mathcal{Y}$. The $O(\sqrt{KT})$ bound has an optimal dependence on $T$; it matches the $\sqrt{T}$ lower bound in the special case of online multiclass classification with bandit feedback (Van der Hoeven et al., 2021, Corollary 1). Furthermore, our bound is better than the existing $O(K\sqrt{T})$ bound of Van der Hoeven (2020) by a factor of $\sqrt{K}$, while it is not directly comparable to the latest $O(K\sqrt{T})$ bound in Van der Hoeven et al. (2021) due to differences in surrogate loss functions. See Appendix C for a more detailed discussion.
@@
While the $O(\sqrt{KT})$ bound is satisfactory when $K=\lvert\mathcal{Y}\rvert$ is small, $K$ can be extremely large in some structured prediction problems: in multilabel classification with $m$ correct labels, we have $K=\binom{d}{m}$, and in ranking problems with $m$ items, we have $K=m!$. To address this issue, we consider a special case of SELF (denoted by SELF* in Table 1), which still includes the aforementioned examples: the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking. A technical challenge to resolve the issue lies in designing an appropriate gradient estimator used in online learning methods. To this end, we draw inspiration from pseudo-inverse estimators used in the adversarial linear/combinatorial bandit literature (Dani et al., 2007; Abernethy et al., 2008; Cesa-Bianchi & Lugosi, 2012). While we cannot naively apply the existing estimators, we design a new gradient estimator that applies to various specific structured prediction problems belonging to the special SELF framework. Armed with this gradient estimator, we achieve a surrogate regret upper bound of $O(T^{2/3})$, which does not explicitly depend on $K$ (Theorem 3.8).
@@
For the delayed feedback setting with a known fixed delay time of $D$, it is actually not difficult to obtain a surrogate regret bound of $O(\sqrt{DT})$ with standard Online Convex Optimization (OCO) algorithms for delayed feedback. Our finding is that we can achieve a surrogate regret bound of $O(D^{2/3}T^{1/3})$ in online structured prediction under delayed full information feedback (Theorems 4.3 and 4.4) by leveraging ODAFTRL (Flaspohler et al., 2021), a Follow-the-Regularized-Leader-type algorithm that achieves an AdaGrad-type regret upper bound in OCO under delayed feedback. This bound is better than $O(\sqrt{DT})$ as $D\leq T$.
@@
Given the contributions so far, it is natural to explore online structured prediction in environments where both delay and bandit feedback are present. We obtain algorithms for this setup by combining the theoretical developments for bandit feedback without delay and delayed full information feedback, offering surrogate regret bounds of $O(\sqrt{DKT})$ (Theorem 5.1) and $O(D^{1/3}T^{2/3})$ (Theorem 5.2).
We validate our algorithms through numerical experiments using both synthetic and real-world data. Specifically, we consider online multiclass classification with bandit feedback. We observe that, depending on the number of classes and the dataset, our algorithm, designed for general structured prediction, can achieve accuracy comparable to existing algorithms specialized for multiclass classification.
@@ 2 Preliminaries
We describe the detailed setup of online structured prediction and key tools used in this work: the Fenchel–Young loss, SELF, and randomized decoding.
@@
2.1 Online Structured Prediction
@@
Here, we describe the problem setting of online structured prediction. Let $\mathcal{X}$ be the input vector space and $\mathcal{Y}$ be the output vector space. Define $K\coloneqq|\mathcal{Y}|$. Following Blondel et al. (2020) and Sakaue et al. (2024), we assume that $\mathcal{Y}$ is embedded into $\mathbb{R}^{d}$ in a standard manner. For example, $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$ in $d$-class multiclass classification.
A linear estimator $\bm{W}\in\mathcal{W}$ for a convex domain $\mathcal{W}$ is used to transform the input vector $\bm{x}$ into the score vector $\bm{W}\bm{x}$. In online structured prediction, at each round $t=1,\dots,T$:
  1. 1.
    The environment selects an input $\bm{x}_{t}\in\mathcal{X}$ and the true output $\bm{y}_{t}\in\mathcal{Y}$;
  2. 2.
    The learner receives $\bm{x}_{t}$ and computes the score vector $\bm{\theta}_{t}=\bm{W}_{t}\bm{x}_{t}$ using the linear estimator $\bm{W}_{t}$;
  3. 3.
    The learner selects a predicted output $\widehat{\bm{y}}_{t}$ based on $\bm{\theta}_{t}$ and incurs a loss of $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$;
  4. 4.
    The learner receives feedback based on the problem setup and updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using an online learning algorithm, $\mathsf{ALG}$.
@@
The goal of the learner is to minimize the cumulative prediction loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$, which is equivalent to minimizing the surrogate regret $\mathcal{R}_{T}$ in (1). We assume that the input and output are generated in an oblivious manner. Note that when $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\dots,\bm{\mathrm{e}}_{d}\}$ and $L(\widehat{\bm{y}}_{t};\bm{y}_{t})=\mathds{1}[\widehat{\bm{y}}_{t}\neq\bm{y}_{t}]$, the above setting reduces to online multiclass classification, which was studied by Van der Hoeven (2020) and Van der Hoeven et al. (2021). We will use $B\coloneqq\mathrm{diam}(\mathcal{W})$, $C_{x}\coloneqq\mathrm{diam}(\mathcal{X})$, and $C_{y}\coloneqq\mathrm{diam}(\mathcal{Y})$ to denote the diameters of the sets $\mathcal{W}$, $\mathcal{X}$, and $\mathcal{Y}$, respectively.
@@
The feedback observed by the learner depends on the problem setting. The most fundamental setting is the full information setup, where the true output $\bm{y}_{t}$ is observed as feedback at the end of each round $t$. This setup was extensively investigated in Sakaue et al. (2024). By contrast, our study investigates the following weaker feedback:
  • Bandit feedback: Only the value of the loss function is observed. Specifically, at the end of each round $t$, the learner observes the target loss value $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$ as feedback.
  • Delayed feedback: The feedback is observed with a certain delay. We consider a fixed $D$-round delay setting, i.e., no feedback is received for round $t\leq D$, and for $t>D$, the learner observes either full information feedback $\bm{y}_{t-D}$ or bandit feedback $L(\widehat{\bm{y}}_{t-D};\bm{y}_{t-D})$.
@@
(1) There exists $\nu>0$ such that for any distinct $\bm{y},\bm{y}^{\prime}\in\mathcal{Y}$, it holds that $\|\bm{y}-\bm{y}^{\prime}\|\geq\nu$. (2) For each $\bm{y}\in\mathcal{Y}$, the target loss function $L(\cdot;\bm{y})$ is defined on $\operatorname{conv}(\mathcal{Y})$, is non-negative, and is affine with respect to its first argument. (3) There exists $\gamma$ such that for any $\bm{y}^{\prime}\in\operatorname{conv}(\mathcal{Y})$ and $\bm{y}\in\mathcal{Y}$, it holds that $L(\bm{y}^{\prime};\bm{y})\leq\gamma\|\bm{y}^{\prime}-\bm{y}\|$ and $L(\bm{y}^{\prime};\bm{y})\leq 1$. (4) It holds that $L(\bm{y}^{\prime};\bm{y})=0$ only if $\bm{y}^{\prime}=\bm{y}$.
As discussed in Sakaue et al. (2024, Section 2.3), these assumptions are natural and hold for a broad range of problem settings and target loss functions, including SELF (see Section 2.4 for the formal definition).

### v2

@@ Abstract
Online structured prediction is a task of sequentially predicting outputs with complex structures based on inputs and past observations, encompassing online classification. Recent studies showed that in the full-information setting, we can achieve finite bounds on the surrogate regret, i.e., the extra target loss relative to the best possible surrogate loss. In practice, however, full-information feedback is often unrealistic as it requires immediate access to the whole structure of complex outputs. Motivated by this, we propose algorithms that work with less demanding feedback, bandit and delayed feedback. For bandit feedback, by using a standard inverse-weighted gradient estimator, we achieve a surrogate regret bound of $O(\sqrt{KT})$ for the time horizon $T$ and the size of the output set $K$. However, $K$ can be extremely large when outputs are highly complex, resulting in an undesirable bound. To address this issue, we propose another algorithm that achieves a surrogate regret bound of $O(T^{2/3})$, which is independent of $K$. This is achieved with a carefully designed pseudo-inverse matrix estimator. Furthermore, we numerically compare the performance of these algorithms, as well as existing ones. Regarding delayed feedback, we provide algorithms and regret analyses that cover various scenarios, including full-information and bandit feedback, as well as fixed and variable delays.
@@ 1 Introduction
In many machine learning problems, given an input vector from a set $\mathcal{X}$ of input vectors, we aim to predict a vector in a finite output space $\mathcal{Y}$. Multiclass classification is one of the simplest examples, while in other cases, output spaces may have more complex structures. Structured prediction refers to such a class of problems with structured output spaces, including multiclass classification, multilabel classification, ranking, and ordinal regression, and it has applications in various fields, ranging from natural language processing to bioinformatics [37, 2]. In structured prediction, training models that directly predict outputs in complex discrete output spaces is typically challenging. Therefore, we often adopt the surrogate loss framework [3]—define an intermediate space of score vectors and train models that estimate score vectors from inputs based on surrogate loss functions. Examples of surrogate losses include squared, logistic, and hinge losses, and a general framework encompassing them is the Fenchel–Young loss [6], which we rely on in this study.
Structured prediction can be naturally extended to the online setting, called Online Structured Prediction (OSP) [35]. In OSP, at each round $t=1,\dots,T$, an environment selects an input–output pair $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{X}\times\mathcal{Y}$. A learner then predicts $\widehat{\bm{y}}_{t}\in\mathcal{Y}$ based on the input $\bm{x}_{t}$ and incurs a loss $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$, where $L\colon\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is the target loss function. Following prior work [38, 39, 35], we focus on the simple yet fundamental case where the learner’s model for estimating score vectors is linear.
Table 1: Upper and lower bounds on the surrogate regret in online multiclass classification and OSP. Here, $T$ is the time horizon, $K=\lvert\mathcal{Y}\rvert$ is the size of the output space, and $D$ is the fixed delay time. Delayed feedback is considered only when “Delayed” appears in the feedback column. In the target loss column, “SELF*” means SELF that satisfies Assumption 3.6. Note that the $O(T^{2/3})$ bounds for SELF* in lines 6 and 9 do not explicitly depend on $K$ but on $d$; in the case of multiclass classification with the 0-1 loss, the dependence on $K$ appears as $d=K$.
Problem setting Feedback Target loss Surrogate regret bound
van der Hoeven et al. [39, Cor. 1] Binary classification Graph bandit 0-1 loss $\Omega(\sqrt{T})$ ($d=2$)
van der Hoeven [38, Thm. 4] Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
van der Hoeven et al. [39, Thm. 1] Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
Sakaue et al. [35, Thms. 7 and 8] Structured prediction Full-info SELF $O(1)$
This work (Theorems 3.5 and D.2) Structured prediction Bandit SELF $O(\sqrt{KT})$
@@
This work (Theorems 4.3 and E.3) Structured prediction Full-info & Delayed SELF $O(\min\{D^{2},D^{2/3}T^{1/3}\})$
This work (Theorem 5.1) Structured prediction Bandit & Delayed SELF $O(\sqrt{(K+D)T})$
@@
The goal of the learner is to minimize the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$. On the other hand, the best the learner can do in the surrogate loss framework is to minimize the cumulative surrogate loss, namely $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$, where $\bm{U}\colon\mathcal{X}\to\mathbb{R}^{d}$ is the best offline linear estimator and $S\colon\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a surrogate loss, which measures the discrepancy between the score vector $\bm{U}\bm{x}_{t}\in\mathbb{R}^{d}$ and $\bm{y}_{t}\in\mathcal{Y}$. Given this, a natural performance measure of the learner’s predictions is the surrogate regret, $\mathcal{R}_{T}$, such that $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}$. The surrogate regret was introduced in the seminal paper by van der Hoeven [38] in the context of online multiclass classification. Recently, Sakaue et al. [35] showed that a finite surrogate regret bound can be achieved for OSP under full-information feedback, i.e., the learner can observe $\bm{y}_{t}$ at the end of each round $t$.
However, the assumption that full-information feedback is available is often demanding, especially when outputs have complex structures. For example, in sequential ad assortment on an advertising platform, we may be able to observe only the click-through rate but not which ads were clicked, which boils down to the bandit feedback setting [24, 19]. Also, we may only have access to feedback from a while ago when designing an ad assortment for a new user—namely, delayed feedback [41, 27]. Similar situations have led to a plethora of studies in various online learning settings. In combinatorial bandits, algorithms under bandit feedback (referred to as full-bandit feedback in their context), instead of full-information feedback, have been widely studied [8, 12, 33, 15]. Delayed feedback is also explored in various other settings, including full-information and bandit feedback [22, 9]. Due to space limitations, we defer a further discussion of the background to Appendix B.
@@
To extend the applicability of OSP, this study develops OSP algorithms that can handle weaker feedback—bandit feedback and delayed feedback—instead of full-information feedback. Following the work of Sakaue et al. [35], we consider the case where target loss functions belong to a class called the Structured Encoding Loss Function (SELF) [10, 5], a general class that includes the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking (see Section 2.4 for the definition). Table 1 summarizes the surrogate regret bounds provided in this study and comparisons with the existing results.
@@
One of the major challenges in the bandit feedback setting is that the true output $\bm{y}_{t}$ is not observed, making it impossible to compute the true gradient of the surrogate loss. To address this, we use an inverse-weighted gradient estimator, a common approach that assigns weights to gradients by the inverse probability of choosing each output, establishing an $O(\sqrt{KT})$ surrogate regret upper bound (Theorems 3.5 and D.2), where $K=\lvert\mathcal{Y}\rvert$ denotes the cardinality of $\mathcal{Y}$. This $O(\sqrt{KT})$ bound has a desirable dependence on $T$, matching an $\Omega(\sqrt{T})$ lower bound known in a problem closely related to online multiclass classification with bandit feedback [39, Corollary 1]. Furthermore, our bound is better than the existing $O(K\sqrt{T})$ bounds [38, 39] by a factor of $\sqrt{K}$, although the bound of van der Hoeven et al. [39] applies to a broader class of surrogate losses and thus it is not directly comparable to ours (see Appendix C for a more detailed discussion). We also conduct numerical experiments on online multiclass classification and find that our methods, which apply to general OSP, are comparable to the existing ones specialized for multiclass classification (see Section H.1).
@@
While the $O(\sqrt{KT})$ bound is satisfactory when $K=\lvert\mathcal{Y}\rvert$ is small, $K$ can be extremely large in some structured prediction problems. In multilabel classification with $m$ correct labels, we have $K=\binom{d}{m}$, and in ranking problems with $m$ items, we have $K=m!$. To address this issue, we consider a special case of SELF (denoted by SELF* in Table 1), which still includes the aforementioned examples: the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking. A technical challenge to resolve the issue lies in designing an appropriate gradient estimator used in online learning methods. To this end, we draw inspiration from pseudo-inverse estimators used in the adversarial linear/combinatorial bandit literature [14, 1, 8]. Indeed, we cannot naively use the existing estimators, and hence we design a new gradient estimator that applies to various structured prediction problems with target losses belonging to the special SELF class. Armed with this gradient estimator, we achieve a surrogate regret upper bound of $O(T^{2/3})$ (Theorem 3.7). This successfully eliminates the explicit dependence on $K$, although the dependence on $T$ increases compared to the $O(\sqrt{KT})$ bound. We also numerically observe the benefit of this approach when $K$ is large, aligning with the implication of the theoretical bounds (see Section H.2).
@@
For the delayed feedback setting, the surrogate regret bounds depend on whether the delay time is fixed or variable. We here describe our results for the known fixed delay time, denoted by $D$, under the full-information setting. It is relatively straightforward to obtain a surrogate regret bound of $O(\sqrt{DT})$ with standard Online Convex Optimization (OCO) algorithms for the delayed feedback. We improve upon this bound and achieve a surrogate regret bound of $O(\min\{D^{2},D^{2/3}T^{1/3}\})$, which is notably independent of $T$ (Theorems 4.3 and E.3). The proof requires carefully bridging different lines of research: an AdaGrad-type OCO algorithm for delayed feedback [17] and surrogate regret analysis in OSP.
@@
Given the contributions so far, it is natural to explore OSP in environments where both delay and bandit feedback are present. We develop algorithms for this setting by combining the theoretical developments for bandit feedback without delay and delayed full-information feedback, achieving surrogate regret bounds of $O(\sqrt{(D+K)T})$ (Theorem 5.1) and $O(D^{1/3}T^{2/3})$ (Theorem 5.2). It is worth noting that similar surrogate regret bounds can also be obtained in the variable-delay setting—where the delay may differ for each round of feedback—for both full-information and bandit feedback. Due to space constraints, most of the details are provided in Appendix G.
@@ 2 Preliminaries
This section describes the detailed setting of OSP and key tools used in this work: the Fenchel–Young loss, SELF, and randomized decoding.
@@
2.1 Online structured prediction (OSP)
@@
We describe the problem setting of OSP. Let $\mathcal{X}$ be the space of input vectors and $\mathcal{Y}$ be the finite set of outputs. Define $K\coloneqq|\mathcal{Y}|$. Following the literature [6, 35], we assume that $\mathcal{Y}$ is embedded into $\mathbb{R}^{d}$ in a standard manner, e.g., $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$ in multiclass classification with $d$ classes.
Let $\mathcal{W}$ be a closed convex domain. A linear estimator $\bm{W}\in\mathcal{W}$ transforms the input vector $\bm{x}$ into the score vector $\bm{W}\bm{x}$. In OSP, at each round $t=1,\dots,T$, the environment selects an input $\bm{x}_{t}\in\mathcal{X}$ and the true output $\bm{y}_{t}\in\mathcal{Y}$. The learner receives $\bm{x}_{t}$ and computes the score vector $\bm{\theta}_{t}=\bm{W}_{t}\bm{x}_{t}$ using the linear estimator $\bm{W}_{t}$. Then, the learner selects a prediction $\widehat{\bm{y}}_{t}$ based on $\bm{\theta}_{t}$, which is called decoding, and incurs a loss of $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$. Finally, the learner receives feedback, which depends on the problem setting, and updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using some online learning algorithm, denoted by $\mathsf{ALG}$, which is applied to the surrogate loss function $\bm{W}\mapsto S(\bm{W}\bm{x}_{t};\bm{y}_{t})$.
@@
The goal of the learner is to minimize the cumulative target loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$, which is equivalent to minimizing the surrogate regret $\mathcal{R}_{T}=\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}-\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}.$ We assume that the input and output are generated in an oblivious manner. Note that when $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\dots,\bm{\mathrm{e}}_{d}\}$ and $L(\widehat{\bm{y}}_{t};\bm{y}_{t})=\mathds{1}[\widehat{\bm{y}}_{t}\neq\bm{y}_{t}]$, the above setting reduces to online multiclass classification, which was studied in prior work [38, 39]. Let $B\coloneqq\mathrm{diam}(\mathcal{W})$ denote the diameter of $\mathcal{W}$ measured by $\|\cdot\|_{\mathrm{F}}$ and $C_{x}\coloneqq\max_{\bm{x}\in\mathcal{X}}\|\bm{x}\|_{2}$ the maximum Euclidean norm of input vectors in $\mathcal{X}$.
@@
The feedback observed by the learner depends on the problem setting. The most basic setting is the full-information setting, where the true output $\bm{y}_{t}$ is observed as feedback at the end of each round $t$, extensively investigated by Sakaue et al. [35]. By contrast, we study the following weaker feedback: In the bandit feedback setting, only the value of the loss function is observed. Specifically, at the end of each round $t$, the learner observes the target loss value $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$ as feedback. In the delayed feedback setting, the feedback is observed with a certain delay. In this paper, we consider especially a fixed $D$-round delay setting, i.e., no feedback is received for round $t\leq D$, and for $t>D$, the learner observes either full-information feedback $\bm{y}_{t-D}$ or bandit feedback $L(\widehat{\bm{y}}_{t-D};\bm{y}_{t-D})$. We also discuss the variable-delay setting in Appendix G.
@@
(1) There exists $\nu>0$ such that for any distinct $\bm{y},\bm{y}^{\prime}\in\mathcal{Y}$, it holds that $\|\bm{y}-\bm{y}^{\prime}\|\geq\nu$. (2) For each $\bm{y}\in\mathcal{Y}$, the target loss function $L(\cdot;\bm{y})$ is defined on $\operatorname{conv}(\mathcal{Y})$, non-negative, and affine w.r.t. its first argument. (3) There exists $\gamma$ such that for any $\bm{y}^{\prime}\in\operatorname{conv}(\mathcal{Y})$ and $\bm{y}\in\mathcal{Y}$, it holds that $L(\bm{y}^{\prime};\bm{y})\leq\gamma\|\bm{y}^{\prime}-\bm{y}\|$ and $L(\bm{y}^{\prime};\bm{y})\leq 1$. (4) It holds that $L(\bm{y}^{\prime};\bm{y})=0$ if and only if $\bm{y}^{\prime}=\bm{y}$.
As discussed in [35, Section 2.3], these assumptions are natural and hold for various structured prediction problems and target loss functions, including SELF (see Section 2.4 for the formal definition).